# A Machine Learning and Finite Element Framework for Inverse Elliptic PDEs via Dirichlet-to-Neumann Mapping

### v1

@@ 4.5 Example 3: Phantom Coefficient with a Zero Source Function
In this subsection, we consider the phantom coefficient, which is used in many medical imaging applications such as EIT [6, 7, 8] and PAT.
In this case, we only consider the fully observed setting, where $\omega=\partial\Omega$. The source function $f$ is set to zero, and the domain $\Omega$ is defined as a circular region:
$\Omega=\{(x,y)\mid(x-0.5)^{2}+(y-0.5)^{2}\leq 0.5^{2}\}.$
The phantom coefficient $k$ within this domain is configured as shown in Figures 15(a) and 20(a). We conduct numerical experiments on two different phantom configurations of $k$, where a slight smoothing effect is applied using a Gaussian kernel. This reduces sharp transitions in the values and yields a more continuous distribution across the domain.
Since the inverse problem is ill-posed, we include a regularization term in the loss function to improve the stability and robustness of the reconstructed coefficient. Specifically, we extend the loss function (3) by introducing a total variation term [6], and use the following loss function:
@@
Here, $\mathfrak{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\partial\Omega}\}$, $\tilde{\mathfrak{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\partial\Omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega)\}$. The positive parameter $\lambda$ acts as a regularization weight, controlling the contribution of the regularization term $\|\nabla\tilde{k}\|_{L^{2}(\Omega)}^{2}$.
The experimental setup for the phantom case is as follows. It is represented by a mesh with $N_{\text{phantom}}=51,433$ points. A Dirichlet boundary condition is imposed, given by $g=\cos(a\pi x)\cos(b\pi y)\big{|}_{\partial\Omega},$ where $a$ and $b$ are uniformly sampled from the interval $[0,2]$.
4.5.1 Example 3-1: Phantom 1
In this case, Phantom 1 is used as $k$ due to its relatively simple shape compared to the other phantom. We set the regularization weight in the loss function (6) as $\lambda=0.001/N_{\text{phantom}}$.
@@ 4.5.1 Example 3-1: Phantom 1
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 12: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 12 compares the results of the model trained with the regularization term included in the loss function (6) and the model trained without it using the loss function (3). The results show no significant difference between the two models, suggesting that the regularization term does not have a major impact due to the relative simplicity of the problem. Additionally, Figure 13 provides a three-dimensional illustration of the phantom configuration.
@@
(b) reconstructed $\tilde{k}$ using (3)
(c) reconstructed $\tilde{k}$ using (6)
Figure 13: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
Figure 14: Relative error of $k$ with and without regularization term over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 14 plots the relative error of $k$ on a logarithmic scale. The results indicate that the model with the regularization term slightly outperforms the one without it, although the difference remains negligible. Both models exhibit a steady decrease in relative error and eventually converge to similar values. This suggests that the regularization term has little impact on the accuracy of the $k$ reconstruction in this case.
4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 1 in Example 3-1 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.2,1.0]$, assuming prior knowledge of this range. Apart from this post-processing step, the experimental setup remains the same as Example 3-1. We set the regularization weight in the loss function (6) as $\lambda=0.01/N_{\text{phantom}}$.
@@ 4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$  (6)
(e) absolute error $|k-\tilde{k}|$
Figure 15: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 15 compares the results of models trained using the loss function with the regularization term (6) and without it (3). The comparison indicates that the inclusion of the regularization term does not significantly affect the final reconstruction, as both models produce similar outcomes. However, a comparison between the results without post-processing (Figure 12) and with post-processing (Figure 15) shows that the region with noticeable error is significantly reduced in the post-processed case. This suggests that the post-processing step plays an important role in improving the spatial accuracy of the reconstructed coefficient.
Figure 16: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 16 presents the relative error in the coefficient $k$ for models trained with and without the regularization term, after applying the post-processing step. As in the case without post-processing, the model trained with the regularization term shows slightly better performance, but the difference remains negligible. This confirms that the regularization term has minimal impact on the accuracy of $k$ reconstruction, even after post-processing.
4.5.3 Example 3-2: Phantom 2
In this case, a more complex phantom, Phantom 2, is used to more clearly assess the effect of the regularization term. We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
@@ 4.5.3 Example 3-2: Phantom 2
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 17: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 17 presents visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6), and their corresponding absolute errors. This comparison illustrates the differences between the models trained with and without the regularization term.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) reconstructed $\tilde{k}$ using (6)
Figure 18: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
To complement this, Figure 18 provides a three-dimensional visualizations of the same results. The overall structure of $k$ is well captured, but the reconstruction tends to overshoot at the peaks, producing sharper and taller spikes than the exact coefficient.
Figure 19: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 19 shows the relative error of $k$ over epochs on a logarithmic scale. While both models show similar convergence during the initial epochs of training, the model without regularization shows a slight increase in error in the later epochs. In contrast, the model with regularization maintains stable convergence throughout, suggesting that the regularization term improves training stability.
4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 2 in Example 3-2 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[1.0,5.0]$, assuming prior knowledge of this range. Apart from this post-processing step, the experimental setup remains the same as Example 3-2. We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 20: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 20 compares the reconstruction results of two models: one trained using the loss function (3), and the other using the loss function (6), which includes a regularization term. Figures 20(c) and 20(e) show the corresponding absolute error distributions. In Figure 20(c), large errors are observed around regions with sharp transitions in $k$, while Figure 20(e) shows smaller errors in these regions. This indicates that the model trained with the regularization term is better at capturing rapid changes and interfaces in the coefficient. These results suggest that the regularization term plays an important role in preserving the structural consistency of $k$, especially in complex phantom configurations.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using the loss function (3) with $\omega=\partial\Omega$
(c) reconstructed $\tilde{k}$ using the loss function (6)
Figure 21: Illustrations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ in three dimensions.
To complement this, Figure 21 provides a three-dimensional visualizations of the same results. The three-dimensional visualizations provide a clearer perspective on the shape and height of the reconstructed peaks. While both models capture the overall structure of $k$, the model with the regularization term—shown in Figure 21(c)—achieves more accurate peak heights and smoother transitions. In contrast, the model without the regularization term, shown in Figure 21(b), shows slight deformation around regions with sharp transitions in $k$.
Figure 22: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 22 shows the relative error of $k$ over epochs, plotted on a logarithmic scale. The model with regularization (red curve) shows a continuously decreasing error, especially after 30 epochs, and finally reaches a significantly lower final error than the model without regularization (blue curve). This behavior confirms that the regularization term not only improves accuracy but also stabilizes the training process, preventing overfitting and helping the model converge more reliably in complex scenarios.
Case Post-Processing MSE (without regularization) MSE (with regularization)
Phantom 1 No 0.00161 0.00155
Yes 0.00031 0.00029
Phantom 2 No 0.26601 0.21033
Yes 0.05494 0.02955
Table 2: MSE for Phantom 1 and Phantom 2 with and without regularization and post-processing.
Table 2 summarizes the MSE for Phantom 1 and Phantom 2 under different combinations of regularization and post-processing. For both phantoms, post-processing significantly improves the reconstruction accuracy, reducing the MSE by up to an order of magnitude. While the inclusion of the regularization term alone has only a marginal effect in Phantom 1, it is more effective in Phantom 2, especially when combined with post-processing. These results suggest that both regularization and post-processing contribute to improving reconstruction accuracy, with their combined effect being particularly important in more complex cases such as Phantom 2.
4.6 Example 4: Discontinuous Coefficient
In this example, we consider a more challenging and interesting case than the previous one, where the coefficient $k$ is discontinuous. The coefficient is defined as
$k(x,y)=\left\{\begin{array}[]{rl}0.9&\text{ if }(x-0.5)^{2}+(y-0.5)^{2}<(0.25)^{2}\\ 0.5&\text{ otherwise}\end{array}\right.$
as illustrated in Figure 23(a).
We conduct experiments for both the fully and partially observed settings. For the fully observed setting, the computational domain, number of mesh points, and Dirichlet boundary condition are the same as in Example 1. The source function is defined as $f(x,y)=k(x,y)(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$. For the partially observed setting, both the Dirichlet boundary condition and the source function are set to zero on the region $\Gamma_{0}$, while all other setups remain the same as in the fully observed setting. In the post-processing step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.4,1.0]$.
In this example, we train the model using the loss function (3), considering both $\omega=\partial\Omega$ and $\omega=\Gamma$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3) with $\omega=\partial\Omega$
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (3) with $\omega=\Gamma$
(e) absolute error $|k-\tilde{k}|$
@@ 4.6 Example 4: Discontinuous Coefficient
Figure 23: Visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ in three dimensions, and the corresponding absolute errors as a heatmap, for two different observation settings. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the error between (a) and (d).
@@
Figure 23 shows the reconstruction results and corresponding absolute errors for both the fully observed ($\omega=\partial\Omega$) and the partially observed ($\omega=\Gamma$) settings. The absolute error maps in Figures 23(c) and 23(e) show that the reconstruction errors are primarily concentrated around regions where $k$ is discontinuous. This suggests that the model struggles to capture sharp transitions, which is expected given the inherent difficulty of learning discontinuous coefficients.
@@
Figure 24: Slices of the exact and reconstructed coefficients along the diagonal line $y=x$.
Figure 24 shows a slice of the reconstructed coefficient $\tilde{k}$ along the diagonal line $y=x$. The blue dashed line represents the reconstruction from the fully observed setting ($\omega=\partial\Omega$), while the red dashed line corresponds to the reconstruction from the partially observed setting ($\omega=\Gamma$). The orange solid line indicates the exact coefficient $k$. Although small errors are observed near the discontinuities, both reconstructions closely approximate the true coefficient with minimal deviation.
@@
Discontinuous Fully $2.28\times 10^{-4}$ $1.72\times 10^{-3}$ 35.51 0.9835
Partially $1.69\times 10^{-4}$ $1.11\times 10^{-3}$ 36.79 0.9896
Table 3: Comparison of MSE, MAE, PSNR, and SSIM for the discontinuous case under fully and partially observed settings.
Table 3 summarizes the MSE for the discontinuous case under both fully and partially observed settings. Similarly to the result for the spatially varying case, the MSE for the partially observed setting is slightly lower than that for the fully observed setting. This may be due to the fact that the reconstruction errors are mainly concentrated near the discontinuities of $k$, and both models achieve sufficiently accurate solutions in general. In such regimes, minor numerical differences during training may result in the partially observed model yielding a slightly better final reconstruction.

### v3

@@ 4.5 Example 3: Phantom Coefficient with a Zero Source Function
In this subsection, we consider the phantom coefficient, which is used in many medical imaging applications such as EIT [Bar21, Cheney99, Patterson05].
In this case, we only consider the fully observed setting, where $\omega=\partial\Omega$. The source function $f$ is set to zero, and the domain $\Omega$ is defined as a circular region: $\Omega=\{(x,y)\mid(x-0.5)^{2}+(y-0.5)^{2}\leq 0.5^{2}\}.$ The phantom coefficient $k$ within this domain is configured as shown in Figures 11(a) and 16(a). We conduct numerical experiments on two different phantom coefficients $k$, where a slight smoothing effect is applied using a Gaussian kernel. This reduces sharp transitions in the values and yields a more continuous distribution across the domain.
To enhance the accuracy in the case where the coefficient varies over the space, we include a regularization term in the loss function to improve the stability and robustness of the reconstructed coefficient [Bar21], and use the following loss function:
@@
Here, $\mathfrak{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\partial\Omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\partial\Omega}\}$, $\tilde{\mathfrak{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\partial\Omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\partial\Omega\}$. The positive parameter $\lambda$ acts as a regularization weight, controlling the contribution of the regularization term $\|\nabla\tilde{k}\|_{L^{2}(\Omega)}^{2}$.
The experimental setup for the phantom case is as follows. The Dirichlet boundary condition is imposed, given by $g(x,y)=\cos(a\pi x)\cos(b\pi y)\big|_{\partial\Omega},$ where $a$ and $b$ are uniformly sampled from the interval $[0,2]$. The circular domain is discretized using a triangular mesh with 51,433 degrees of freedom for the linear Lagrange finite element space. This experiment uses a Cauchy dataset containing 2,048 samples. The learning rate is set to 0.001 for the first 50 epochs and reduced to 0.0001 for the remaining 50 epochs.
4.5.1 Example 3-1-1: Phantom 1
In this case, Phantom 1 (Figure 11(a)) is used as the coefficient $k$ due to its relatively simple structure compared to the other phantom. The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-8}$.
(a) Phantom 1: exact $k$
(b) $\tilde{k}$ using (2)
(c) absolute error |kk~||k-$\tilde${k}|
(d) $\tilde{k}$ using (5)
(e) absolute error |kk~||k-$\tilde${k}|
Figure 11: Example 3-1-1. Phantom 1. Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 11 compares the results of two models: one trained with the regularization term included in the loss function (5), and the other trained without it using the loss function (2). The results indicate that the inclusion of the regularization term does not lead to a significant difference in performance. Additionally, Figure 12 provides three-dimensional visualizations of the phantom coefficient.
@@ 4.5.1 Example 3-1-1: Phantom 1
(b) $\tilde{k}$ using (2)
(c) $\tilde{k}$ using (5)
Figure 12: Example 3-1-1. Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5).
Figure 13: Example 3-1-1. Relative error of $\tilde{k}$ with and without regularization over epochs (log scale). Each point represents the averaged value over a 10-epoch interval.
Figure 13 shows the relative error of $\tilde{k}$ on a logarithmic scale. The results indicate that the model trained with the regularization slightly outperforms the model trained without regularization, although the difference remains negligible. Both models exhibit a steady decrease in relative error and eventually converge to similar values. This suggests that, in this case, the regularization has little impact on the reconstruction accuracy of the coefficient $k$.
4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 1 in Example 3-1-1 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[0.2,1.0]$, based on the assumed prior knowledge of the range of $k$. The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-7}$. All other experimental configurations remain the same as in Example 3-1-1.
(a) Phantom 1: exact $k$
(b) $\tilde{k}$ using (2)
(c) absolute error |kk~||k-$\tilde${k}|
(d) $\tilde{k}$ using (5)
(e) absolute error |kk~||k-$\tilde${k}|
Figure 14: Example 3-1-2. Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions with the proposed post-processing. (c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 14 compares the results of models trained with and without the regularization term, using the loss functions (5) and (2), respectively. The comparison indicates that the inclusion of the regularization term does not significantly affect the final reconstruction, as both models produce similar outcomes. In contrast, a comparison between the results without post-processing (Figure 11) and with post-processing (Figure 14) shows a significant reduction in the region with noticeable error after post-processing. This suggests that the post-processing step plays an important role in improving the spatial accuracy of the reconstructed coefficient.
Figure 15: Example 3-1-2. Relative error of $\tilde{k}$ with and without regularization terms, using the post-processing, over epochs (log scale). Each point represents the average over a 10-epoch interval.
Figure 15 presents the relative error in the coefficient $k$ for models trained with and without the regularization term, after applying the post-processing step. As in the case without post-processing, the model trained with the regularization term shows slightly better performance, although the difference remains negligible.
4.5.3 Example 3-2-1: Phantom 2
In this case, a more complex phantom, Phantom 2 (Figure 16(a)), is used to more clearly assess the effect of the regularization term. The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-6}$.
(a) Phantom 2: exact $k$
(b) $\tilde{k}$ using (2)
(c) absolute error |kk~||k-$\tilde${k}|
(d) $\tilde{k}$ using (5)
(e) absolute error |kk~||k-$\tilde${k}|
Figure 16: Example 3-2-1. Phantom 2. Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 16 presents visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5), and their corresponding absolute errors.
@@ 4.5.3 Example 3-2-1: Phantom 2
(b) $\tilde{k}$ using (2)
(c) $\tilde{k}$ using (5)
Figure 17: Example 3-2-1. Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5).
Furthermore, Figure 17 provides three-dimensional visualizations of the same results. The overall structure of $k$ is well captured, but the reconstruction tends to overshoot the peaks, producing sharper and taller spikes than the exact coefficient.
Figure 18: Example 3-2-1. Relative error of $\tilde{k}$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10-epoch interval.
Figure 18 shows the relative error of $\tilde{k}$ over epochs on a logarithmic scale. While both models show similar convergence in the initial training epochs, the model trained without regularization shows a slight increase in error in the later stages. In contrast, the model with regularization maintains stable convergence throughout, suggesting that the regularization term contributes to improved training stability.
4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 2 in Example 3-2-1 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[1.0,5.0]$, based on the assumed prior knowledge of the range of $k$. The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-6}$. All other experimental configurations remain the same as in Example 3-2-1.
(a) Phantom 2: exact $k$
(b) $\tilde{k}$ using (2)
(c) absolute error |kk~||k-$\tilde${k}|
(d) $\tilde{k}$ using (5)
(e) absolute error |kk~||k-$\tilde${k}|
Figure 19: Example 3-2-2.Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions with post-processing. (c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 19 compares the reconstruction results of two models: one trained using the loss function (2), and the other using the loss function (5), which includes a regularization term. Figures 19(c) and 19(e) show the corresponding absolute error distributions. In Figure 19(c), large errors are observed near regions with sharp transitions in $k$, while Figure 19(e) shows smaller errors in these regions. This indicates that the model trained with the regularization term is more effective at capturing rapid variations and interfaces in the coefficient. These results suggest that the regularization term plays a critical role in preserving the structural consistency of $k$, particularly in complex phantom coefficients.
@@ 4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
(b) $\tilde{k}$ using (2)
(c) $\tilde{k}$ using (5)
Figure 20: Example 3-2-2. Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5) with post-processing.
To complement this, Figure 20 presents three-dimensional visualizations of the same results. These visualizations provide a clearer perspective on the shape and height of the reconstructed peaks. While both models capture the overall structure of $k$, the model with the regularization term—shown in Figure 20(c)—achieves more accurate peak heights and smoother transitions. In contrast, the model without the regularization term, shown in Figure 20(b), shows slight deformation near regions with sharp transitions in $k$.
Figure 21: Example 3-2-2. Relative error of $\tilde{k}$ with and without regularization terms, using the post-processing, over epochs (log scale). Each point represents the averaged value over a 10-epoch interval.
Figure 21 shows the relative error of $\tilde{k}$ over epochs, plotted on a logarithmic scale. The model with regularization (red curve) shows a consistently decreasing error, especially after 30 epochs, and ultimately reaches a significantly lower final error than the model without regularization (blue curve). This behavior suggests that the regularization term not only improves reconstruction accuracy but also enhances training stability by promoting reliable convergence in complex scenarios.
Case Post-Processing Regularization MSE MAE PSNR SSIM
Phantom 1 No No 0.00162 0.02349 28.17 0.9575
Yes 0.00155 0.02291 28.33 0.9583
Yes No 0.00031 0.00721 35.03 0.9868
Yes 0.00030 0.00708 35.30 0.9872
Phantom 2 No No 0.26602 0.25938 23.66 0.9336
Yes 0.21033 0.23995 24.20 0.9355
Yes No 0.05494 0.11752 26.58 0.9488
Yes 0.02955 0.09589 29.27 0.9653
Table 2: Example 3.1-1 - 3.2-2. Comparison of metrics—MSE, MAE, PSNR, and SSIM—for Phantom 1 and Phantom 2 under different settings of post-processing and regularization.
Table 2 summarizes the performance metrics—MSE, MAE, PSNR, and SSIM—for Phantom 1 and Phantom 2 under different combinations of post-processing and regularization. For both phantoms, post-processing leads to significant improvements in all metrics. These results suggest that while both regularization and post-processing improve reconstruction accuracy, their combined effect becomes particularly beneficial in more complex scenarios, such as Phantom 2.
4.6 Example 4: Discontinuous Coefficient
In this final example, we consider a more challenging and interesting case than the previous one, where the coefficient $k$ is discontinuous. The coefficient is defined as
$k(x,y)=\left\{\begin{array}[]{rl}0.9&\text{ if }(x-0.5)^{2}+(y-0.5)^{2}<(0.25)^{2},\\ 0.5&\text{ otherwise},\end{array}\right$.
as illustrated in Figure 22(a).
We conduct experiments for both the fully and partially observed settings. All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-2. The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$. For the partially observed setting, this source function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$. We also include the post-processing step by applying the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[0.4,1.0]$, based on the assumed prior knowledge of the range of $k$. In this example, we train the model using the loss function (2), considering both $\omega=\partial\Omega$ and $\omega=\Gamma$.
@@ 4.6 Example 4: Discontinuous Coefficient
(b) $\tilde{k}$ using (2) with $\omega=\partial\Omega$
(c) absolute error |kk~||k-$\tilde${k}|
(d) $\tilde{k}$ using (2) with $\omega=\Gamma$
(e) absolute error |kk~||k-$\tilde${k}|
@@
Figure 22: Example 4. Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the corresponding absolute errors as a heatmap, for two different observation settings. (c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the error between (a) and (d).
@@
Figure 22 shows the reconstruction results and the corresponding absolute errors for both the fully observed and partially observed settings. The absolute error maps in Figures 22(c) and 22(e) show that the reconstruction errors are primarily concentrated near regions where $k$ is discontinuous. This indicates that the model has difficulty capturing sharp transitions, which is expected due to the inherent challenge of learning discontinuous coefficients.
@@
Figure 23: Example 4. An illustration of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ under fully and partially observed settings, plotted along the diagonal line $y=x$.
Figure 23 shows a slice of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ under the fully and partially observed settings along the diagonal line $y=x$. The solid orange line represents the exact coefficient, while the blue and red dashed lines correspond to the reconstructions from the fully and partially observed settings, respectively. Although slight differences appear near the discontinuities, both reconstructions closely follow the exact coefficient with minimal deviation.
@@
Discontinuous Fully $4.36\times 10^{-4}$ $2.81\times 10^{-3}$ 32.69 0.9733
Partially $5.02\times 10^{-4}$ $3.30\times 10^{-3}$ 32.08 0.9676
Table 3: Example 4. Comparison of MSE, MAE, PSNR, and SSIM for the discontinuous case under fully and partially observed settings.
Table 3 presents the performance metrics—MSE, MAE, PSNR, and SSIM—for the discontinuous case under both the fully and partially observed settings. All metrics show slightly better performance in the fully observed setting, indicating a more accurate reconstruction overall. However, the partially observed setting also achieves highly competitive results. These findings suggest that the proposed method remains robust even with limited boundary data, particularly since reconstruction errors are primarily localized near discontinuities.